Hello tudor,
A friend ask me this and so i started helping them. However, i got 160 lb. However, it sounded a bit weird so i want to double check.
How many pounds of seed worth $1.05 per pound must be mixed with 30 pounds of seed worth $.90 per pound in order to produce a mixture to sell for $1.00 per pound?

Answer by Earlsdon(6294)

Let x = the number of lbs of seed worth $1.05 per pound.
x($1.05)+30($0.90) = (x+30)($1.00) Simplify and solve for x.
1.05x+27 = x+30 Subtract x from both sides.
0.05x+27 = 30 Subtract 27 from both sides.
0.05x = 3 Finally, divide both sides by 0.05
x = 60lbs.
P.S. I'm not a Tudor nor a Plantaganet.